OpenStudy (erinweeks):

Find the products AB and BA to determine whether B is the multiplicative inverse of A.

OpenStudy (erinweeks):

A=

OpenStudy (erinweeks):

B=

OpenStudy (erinweeks):

A. B = A-1 B. B ≠ A-1

OpenStudy (mandre):

Do you know how to do matrix multiplication?

OpenStudy (mandre):

And do you know what the answer must be for B to be A's multiplicative inverse?

OpenStudy (erinweeks):

a little bit not good at all.

OpenStudy (erinweeks):

& no i dont know what the answer must be

OpenStudy (mandre):

\[\left[\begin{matrix}A11&A12&A13\\A21&A22&A23 \\ A31&A32 &A33\end{matrix}\right].\left[\begin{matrix}B11&B12&B13\\B21&B22&B23 \\ B31&B32 &B33\end{matrix}\right]=\left[\begin{matrix}C11&C12&C13\\C21&C22&C23 \\ C31&C32 &C33\end{matrix}\right]\] A11 is matrix A row 1, column 1, A22 is matrix A row 2 column 2. You with me?

OpenStudy (erinweeks):

yea so far

OpenStudy (mandre):

C11 = A11*B11 + A12*B21 + A13*B31 That's the sum of the products of Row 1 of matrix A and Column 1 of B to give you C11. C12 = A11*B12 + A12*B22 + A13*B32 That's the sum of the products of Row 1 of matrix A and Column 2 of B to give you C12. For each member of C you just take not it's position then you sum the products of it's equivalent Row in A and the equivalent column in B. C32 = Row 3 of A and Column 2 of B.

OpenStudy (mandre):

That should be: For each member of C you just take NOTE OF it's position then you sum the products of it's equivalent Row in A and the equivalent column in B.

OpenStudy (erinweeks):

okay now im getting a little confused

OpenStudy (mandre):

\[\left[\begin{matrix}.&.&.\\A21&A22&A23 \\ .&. &.\end{matrix}\right].\left[\begin{matrix}.&.&B13\\.&.&B23 \\ .&. &B33\end{matrix}\right]=\left[\begin{matrix}.&.&.\\.&.&C23 \\ .&. &.\end{matrix}\right]\] C23 = A21*B13 + A22*B23 + A23*B33 I dunno if that makes it more clear.

OpenStudy (mandre):

It can get confusing I know. Took me a while to get this. Each member of C must be done the same way.

OpenStudy (erinweeks):

yea it does but were not doing my problem.

OpenStudy (mandre):

Yes we are What is 1*1 + 0*(-1) + 0*0? I'm trying to show you how to do it. It's difficult to explain using just 0s and 1s.

OpenStudy (erinweeks):

oh lol .. its okay but 1!

OpenStudy (mandre):

I'm hoping you can do it yourself now as I need to get back to work. Sorry.

OpenStudy (erinweeks):

nah ill just get the help tomo